OpenStudy (anonymous):

Matthew currently has enough money to buy 45 toy cars. If the cost of each car was 10 cents less, Matthew could buy 5 more cars. How much money does Matthew have to spend on toy cars?

OpenStudy (whpalmer4):

let \(p\) be the price of a car, in cents. He has enough to buy 45 cars, so he has 45p. Now to nail down the value of p, we also know that 45p = (45+5)(p-10), because he can afford 5 more cars if each one costs 10 cents less, or p-10. \[45p = (45+5)(p-10)\]\[45p = 50(p-10)\]

OpenStudy (whpalmer4):

Use the distributive property on that, then solve for \(p\). When you have \(p\), multiply it by 45 to find out how much money he has.

OpenStudy (cruffo):

Keep in mind that your answer will be in cents, not dollars :)